OpenStudy (anonymous):

(a) The inverse function of f(x)=log(x) is f−1(x)= 1/log(x) true false (b) If a and b are positive constants, then y=ln(ax+b) has no vertical asymptote. true false

OpenStudy (ash2326):

@nohegarcia if f(x) is an invertible function, with inverse function \(f^{-1}(x)\) then \[f(f^{-1} ((x))=f^{-1}( f(x))=x\] check if this is satisfied in the first question

OpenStudy (ash2326):

@nohegarcia do you get my point?

OpenStudy (anonymous):

i do i judt dont know how to get there

OpenStudy (anonymous):

nevermind i found it thanks!

OpenStudy (ash2326):

We have \[f(x)=\log (x)\] \[f^{-1} (x)=1/\log x\] \[f(f^{-1} (x))=\log (\frac 1 {\log x})\ne x\] and \[f^{-1}(f(x))=\frac{1}{\log ({\log x}})\ne x\] so is this true or false?

OpenStudy (anonymous):

false cause 1/log(x) is reciprocal not inverse

OpenStudy (ash2326):

good:D